OpenStudy (anonymous):

Find the sum of the first 12 terms of the sequence 1, -4, -9, -14, . . .

OpenStudy (anonymous):

okay I'll show you a cool way of doing this

OpenStudy (anonymous):

use the equation a+(n-1)d where a= first term (1 here) d= difference between two elements (which is -5 here)

OpenStudy (anonymous):

so we get 6-5n

OpenStudy (anonymous):

and name that value as x=6-5n

OpenStudy (anonymous):

to get the sum, use the equation s=n(a+x)/2

OpenStudy (anonymous):

where n=12 a=1 and x=6-5(12)

OpenStudy (anonymous):

hope this helped

OpenStudy (anonymous):

It did, thank you!

OpenStudy (anonymous):

you are welcome